OpenStudy (anonymous):

m+n/5 m=23 and n=7 m+n/5= ????

OpenStudy (anonymous):

this is how i solved it... 23+7/5 23+7=30 5/30=6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

saweet!!

OpenStudy (anonymous):

m=23 and n=7 23+7/5= 23+(7)/(5) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 5. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. 23*(5)/(5)+(7)/(5) Complete the multiplication to produce a denominator of 5 in each expression. (115)/(5)+(7)/(5) Combine the numerators of all fractions that have common denominators. (115+7)/(5) Add 7 to 115 to get 122. (122)/(5)

OpenStudy (anonymous):

this is what I got

OpenStudy (anonymous):